(4 points) A 13-year 3.4% bond with a face value of $180,000 is issued on November 21, 2020. It can be redeemed (at the discretion of the issuer) on any coupon date in 2028, 2029, 2030 or on maturity on November 21, 2033.
a) Find how much an investor should pay for the bond on the date it is issued in order to have a minimum yield of 3%
1
Expert's answer
2020-12-10T13:42:36-0500

Coupon:

"180,000*0.034=6120"

Price:

"\frac{180,000}{1.03^{13}}+6,120*\frac{\frac{1}{1.03}*(\frac{1}{1.03}^{13}-1)}{\frac{1}{1.03}-1}=187,657.17"